# Pyramid-Monozone Synergistic Grasping Policy in Dense Clutter

### v1

@@ IV-A1 Top View Alignment
In this section, we first align the pixel corresponding to the minimum depth value within the depth camera view (e.g., 1280 $\times$ 720) with the center pixel of this view by moving the robot, we denote the view before and after movement as $V$ and $V^{\prime}$. After the global view alignment, to assist the grasping model analyze local information and minimize the influence of depth camera errors and noises, we refine the alignment process by narrowing the view to 224 $\times$ 224 and performing two additional consecutive alignment operations, the view after this movement is denoted as $V^{\prime\prime}$. The choice of 224 $\times$ 224 is because grasping models typically accept images of this size as input. Finally, we use the Segment Anything Model (SAM) [21] to segment the aligned view to obtain the mask $M_{f}$, focusing only on the object corresponding to the minimum depth value. That is, the center pixel $V^{\prime\prime}(c)$ of $V^{\prime\prime}$ serves as the prompt for this segmentation.
Through this series of steps, PSP not only effectively segments the topmost object in a densely cluttered scene, which consistently appears near the center of the camera view, but also mitigates the issue of partially visible graspable objects caused by input size constraints of the grasping model. It is important to note that this method is not multi-view grasping [11], which is a complex optimization process. In contrast, our method is more direct and efficient.
IV-A2 Cross-prompted Segmentation
Here, we propose the cross-prompt to optimize segmentation in PSP because the single-point prompt is highly unstable in densely cluttered environments. And, this issue is particularly pronounced when the object’s appearance is complex, such as the food packaging, where only part of the object is segmented. So, this limitation adversely impacts subsequent prediction and sampling of grasp candidates.
Specifically, it begins by applying the Sobel operator [22] to extract the edge of the mask $M_{f}$ obtained from the initial single-point prompt segmentation. We then search the two most distant pixels, which we refer to as ${p_{m}}$ and ${p_{m}^{\prime}}$ within this edge. As shown in Eq. 4, $P_{e}$ means the set of all pixels on this edge, and ${P_{e}\times P_{e}}$ represents the Cartesian product of the set $P_{e}$. Next, we calculate the perpendicular line $\perp_{(p_{m},p_{m}^{\prime})}$ connecting ${p_{m}}$ and ${p_{m}^{\prime}}$, and intersecting this perpendicular line with the edges yields another pair of the most distant pixels, which we refer to as ${p_{m_{p}}}$ and ${p_{m_{p}}^{\prime}}$, as shown in Eq. 5 and Fig. 3. (a). These four points are then used as prompts to perform the second segmentation, resulting in ${M_{s}}$. Finally, ${M_{s}}$ is refined by image dilation processing: a depth threshold is applied and pixels from the first prompt serve as initial points for segmentation to produce ${M_{d}}$. By combining ${M_{d}}$ and ${M_{p}}$, we obtain the refined ${M_{r}}$, as shown in Algorithm 1. This method effectively exploits the geometric constraints of the initial mask, thus achieving significant results in the second segmentation, as demonstrated in our ablation studies.
$(p_{m_{p}},p_{m_{p}}^{\prime})=\perp_{(p_{m},p_{m}^{\prime})}\cap{P_{e}}$ (5)
IV-B Monozone Sampling Policy (MSP)
After isolating the topmost object by PSP, we will sample the grasp candidates within this object. Here, we discard quality score-based sampling and propose the Monozone Sampling Policy (MSP) to sample grasp candidates for this object. In other words, we do not use the instance mask to guide the quality score-based sampling, but instead focus directly on sampling these grasp candidates within the mask. Moreover, MSP comprehensively considers all parameters of the grasp candidates (rotation $\theta$, position $(x,y)$, and width $w$) and analyzes the relationship between the grasp candidates, the object being grasped, and the neighboring objects of the object being grasped. Next, we will elaborate on these components.
IV-B1 Grasp Angle Calibration
An inappropriate grasp angle can easily cause the object to slip or fall during grasping due to the uneven force distribution on both sides of the parallel-jaw gripper. Therefore, ensuring that all grasp candidates (denoted as $G$) are at the optimal angle before sampling is essential, as it directly determines the subsequent sampling effectiveness. Here, we use the grasping model in [13] to obtain the grasp candidates. Specifically, we first extract the edge of the instance mask. For each grasp candidate, we rotate them clockwise in 2 degree intervals until they reach 360 degrees. For each rotation $r$, we find four intersection points between the two long sides of the grasp candidate and this edge, that is, $p_{t_{l}}$, $p_{b_{l}}$, $p_{t_{r}}$, and $p_{b_{r}}$.
@@ IV-B1 Grasp Angle Calibration
Figure 3: The visualization of cross-prompted segmentation (a) and grasp angle calibration (b).
@@
2:  Output: Refined mask $M_{r}$// Top View Alignment: First conduct global view alignment to get to ${V^{\prime}}$, then conduct local view alignment to get ${V^{\prime\prime}}$, finally segment the object in ${V^{\prime\prime}}$ to get ${M_{f}}$.
3:  ${V^{\prime}}\leftarrow V$, $V^{\prime\prime}\leftarrow V^{\prime}$
4:  ${M_{f}}\leftarrow SAM(V^{\prime\prime}(c))$// Cross-prompted Segmentation: First find four points, then use them to perform segmentation to generate ${M_{s}}$, finally refine ${M_{s}}$ to obtain ${M_{r}}$ based on ${M_{d}}$.
@@
Subsequently, we calculate the angle $\theta^{\prime}$ between the vector $\mathbf{v}_{p_{l}}$ determined by $p_{t_{l}}$ and $p_{b_{l}}$ and the vector $\mathbf{v}_{g_{u}}$ of the long upper side of this grasp candidate, and similarly to get the angle $\theta^{\prime\prime}$ between the vector $\mathbf{v}_{p_{r}}$ determined by $p_{t_{r}}$ and $p_{b_{r}}$ and the vector $\mathbf{v}_{g_{u}}$, shown in Fig. 3. (b). By subtracting 90 degrees from each of these angles, taking the absolute value, and summing them, we obtain the angle difference for each rotation. Finally, we select the rotation with the smallest angle difference ${R}^{*}$(in Eq. 6) and use this rotated box as the new grasp candidate.
[…]
IV-B2 Monozone Grasp Analysis
After grasp angle calibration, we individually analyze each grasp candidate $g_{i}$ to determine the optimal grasp. Specifically, denoting $G^{\prime}$ as grasp candidate sets after angle calibration, we first examine the relationship between each grasp candidate $g_{i}$ and the instance mask $M_{r}$ by checking whether the pixels ${P_{i}}^{s}$ along the two short sides of the grasp candidate fall within the instance mask. So, $g_{i}$ will be filtered out if any pixel $p$ in ${P_{i}}^{s}$ is within mask $M_{r}$. Additionally, to ensure that grasp candidates are concentrated within $M_{r}$, we filter out $g_{i}$ whose centers $c_{i}$ are not within $M_{r}$ and will get grasp candidate sets $G^{\prime\prime}$, this is shown in Eq. 7. However, these processes probable filter all $g_{i}$ out, we adaptively rotate the viewpoint clockwise to alter grasp candidate. That is, if no $g_{i}$ are available from the current viewpoint, we rotate the image 30 degrees at a time and repeat the previous steps until the available grasp candidate are found, as shown in Fig. 4.
Next, based on $G^{\prime\prime}$, we analyze their relationship with adjacent objects by setting a depth threshold $T_{d}$, that is, if the depth difference between any $p$ and $c_{i}$ exceeds $T_{d}$, the grasp candidate $g_{i}$ will be filtered out and get grasp candidate sets $G^{\prime\prime\prime}$, as shown in Eq. 8. So the above steps can effectively filter out most grasp candidates $g_{i}$ that could lead to self-collisions, collisions with adjacent objects, and collisions with the ground.
Finally, based on the sampled grasp candidate sets $G^{\prime\prime\prime}$, we use our previous method [23] to select the $g_{i}$ with the smallest center pixel depth value $d(c_{i})$ as the optimal grasp $g^{*}$, which is shown in Eq. 9. This operation ensures a more stable and safer grasp, especially when dealing with the object with complex shapes.
IV-B3 Optimal Grasp Refinement
Although the optimal grasp $g^{*}$ was obtained through the previous coarse-to-fine analysis, this analysis will be influenced by random errors from the depth camera, where $g^{*}$ might still result in collisions with adjacent objects during grasping execution. Spurred from [24], which defines a series of intervals within the grasp box and adjusts the grasp width and position based on the relationships between these intervals. However, this method relies on local regions of the depth image with errors, and it is computationally cumbersome. So, unlike this method, we directly find the minimum rectangle $R_{ec}$ intersecting the optimal grasp $g^{*}$ and the instance mask $M_{r}$, and this process can be represented by $f(g^{*},M_{r})$. Then calculate the shortest width $w_{s}$ and new center point $c^{\prime}_{i}$ of this grasp through ${min{R_{ec}}(w)}$ and ${R_{ec}}(c)$. Additionally, to mitigate the impact of hand-eye calibration errors, we further expand $w_{s}$ to $w^{\prime}_{s}$ by adding some of the hand-eye calibration translation errors $e_{c}$ in the $X$, $Y$-axes. Finally, we use $w^{\prime}_{s}$ and $c^{\prime}_{i}$ as the new width and center of the grasp for optimal grasp, and denote the final grasp as $g^{*}_{f}$. As shown in Eq. 10, Eq. 11, and Algorithm 2.
$R_{ec}=f(g^{*},M_{r})$ (10)
Figure 4: The visualization of adaptive viewpoint rotation in MGA
@@ IV-B3 Optimal Grasp Refinement
6:  end for// Monozone Grasp Analysis: Filter out $g_{i}$ that results in self-collision, collision with adjacent objects, or collision with the ground to obtain the optimal grasp $g^{*}$.
@@ V Experiments
In this section, we validate the effectiveness of our method by conducting extensive experiments. Firstly, we compare it with baseline grasping methods in various mid-clutter (up to 20 objects) and high-clutter scenes(up to 50 objects). Then we extend the cluttered objects to 100 and analyze the effectiveness of PSP and MSP.
@@ V-A1 Setting for Grasping Model
The baseline methods we will compare are categorized into two groups. The first group includes GGCNN [10], GGCNN2 [11], GRconvnet [13], SEnet [12], and FCGnet [14], which are suitable for mid-clutter scenarios. The second group comprises DexNet 4.0 [4] and GraspNet [5], which are tailored for high-clutter scenarios. For the first group, since the pre-trained models were all trained on the Cornell Grasping Dataset [8], their performance in cluttered environments is limited. Therefore, we merge the OCID Grasping Dataset [6] with the cluttered scene into the Cornell Grasping Dataset and retrain these models using the parameter settings specified in their original papers (except, all using the RGBD modality). Specifically, we randomly shuffle the entire dataset, using 90% for training and 10% for testing. During training, the data are uniformly cropped to fit the acceptable sizes, the number of training epochs is set to 50, and data augmentation (random zoom and random rotation) is applied. For testing, we use the same metric [11] to report the detection accuracy (Acc) of these methods.
@@
V-A2 Setting for Real Grasping
Our overall grasping system consists primarily of an Intel RealSense D435 depth camera and an Xarm5 robot. We employ an eye-in-hand grasping architecture, with the camera mounted on the robot’s end effector and facing downward. The objects used in our grasping experiments, including 300 novel objects: 100 household objects, 100 snacks, 50 ragdolls, and 50 toys and the grasping difficulty increases successively, as shown in Fig. 5.
Before grasping, we use the method from [10] to fill the depth hole and set a depth value threshold (with the upper limit close to the grasping ground and the lower limit at 10 cm to the camera) to ensure that the grasp is executed within this range. During grasping, each method is tested in five trials per experiment, and the number of failed grasps in each trial (T) is recorded. The grasp success rate (GSR) is calculated by dividing the total number of successful grasps by the total number of grasp attempts across five trials. In addition, to improve experimental safety and ensure all objects are grasped in each trial, we provide minimal manual assistance during the experiments. Specifically, if an object fails to be grasped or recognized 2-3 times, we manually pick up the object and count it as a failure. Additionally, if an object moves out of the camera view, it is repositioned with manual intervention. Similarly, if a grasped object moves out of the robot’s range, causing it to stop, the object is repositioned manually too, and the code will be re-executed.
V-B Comparision Studies
V-B1 Comparision with Baseline Methods in Mid-clutter
In this section, we compare PMSGP with the baseline methods in the first group. We used 10 snacks and 10 household objects to form a mid-clutter scene. The results are shown in I. PMSGP achieves a GSR of 93.5% (100/107), with only 7 total grasp failures, which is far superior to other baselines, except for FCGnet. Additionally, we find that some baselines perform well on the benchmark but not in real grasping. For example, GGCNN2 has a GSR of only 47.6% (100/210) with a total of 110 grasp failures, indicating that this method does not generalize well to novel objects in mid-clutter.
Figure 5: The objects for the grasping experiment. On the first row, the left side are toys, and the right side are ragdolls. On the second row, the left side are snacks, and the right side are household objects.
TABLE I: The GSR between PMSGP and first-group baseline methods
@@ V-B1 Comparision with Baseline Methods in Mid-clutter
Figure 6: The grasping performance of PMSGP on objects of different grasping difficulties at different scene complexity. The first row is a mid-clutter scene consisting of 20 ragdolls, the second row is a high-clutter scene consisting of 50 snacks, and the last row is an extreme-clutter scene consisting of 100 household objects. The scene complexity and object grasping difficulty of each row increase in sequence.
TABLE II: The GSR between PMSGP and second-group baseline methods
Methods Ragdolls Snacks Toys Households

### v2

@@ IV-A1 Top View Alignment
We first define the robot’s workspace (the robot base coordinate) with the $X$ and $Y$ axis limited by the edges of a 120 $cm$$\times$ 80 $cm$ table. The range of the $Z$-axis is limited by the maximum distance to the tabletop (40 $cm$) and the minimum distance (10 $cm$) to prevent collisions between the gripper fingers and the table. The camera is mounted at the distal end of the robot arm to which the gripper (about 10 $cm$) is also removably mounted. Before each grasping attempt, we set the robot to a pre-specified position (40 $cm$ above the center of the table) and ensure that the camera covers the entire pile of objects on the table. Then, we conduct the global view alignment based on the initial depth image $V$ and its center $V(c)$ of the camera. Specifically, we align the pixel corresponding to the minimum depth value (among 1280 $\times$ 720 pixels) in $V$ with $V(c)$ by moving the camera mounted on the robot. We denote the depth image after the camera movement by $V^{\prime}$. In theory, a one-time global view alignment should suffice. However, in practice, due to errors introduced by the depth camera (both intrinsic errors and those caused by exposure), the pixel location corresponding to the minimum depth value used for the aforementioned global view alignment may differ from the true location.
Therefore, we narrow the region of interest and perform local view alignment using the minimum depth value pixel in the region to mitigate the impact of depth errors and ensure that the aligned object sits also on the top of the regional pile. Specifically, to improve efficiency, we use the central block with 224 $\times$ 224 pixels of the original 1280 $\times$ 720 image to perform two consecutive local alignments, in the same way as in the global view alignment. Let $V^{\prime\prime}$ and $V^{\prime\prime\prime}$ denote the depth images after each local alignment operation, respectively. One reason for choosing the 224 $\times$ 224 image block is that grasp detection models typically accept images of this size as input. The sequence of alignment is illustrated in Fig. 3. We assume that all minimum depth value pixels stay in the same object with the camera position alignments. Finally, we use the Segment Anything Model (SAM) [30] to segment the $V^{\prime\prime\prime}$ to obtain the segmented RGB image with mask $M_{f}$. The segmentation extracts the object image that contains the pixel of the minimum depth value. Intuitively, the center pixel $V^{\prime\prime\prime}(c)$ of $V^{\prime\prime\prime}$ serves as a prompt for the segmentation.
Figure 3: Sequence of top view alignment: (a) global view alignment, (b) first local view alignment, (c) second local view alignment, and (d) final aligned view. $V(c)$, $V^{\prime}(c)$, $V^{\prime\prime}(c)$, and $V^{\prime\prime\prime}(c)$ are denoted by red point, green point, blue point, and purple point.
@@
Figure 4: Visualization of (a) cross-prompted segmentation and (b) grasp angle calibration. In (a), $p_{m_{p}}$ and $p_{m_{p}}^{\prime}$ are denoted by orange points, $p_{m}$ and $p_{m}^{\prime}$ by green point. In (b), $p_{t_{l}}$, $p_{b_{l}}$, $p_{t_{r}}$, and $p_{b_{r}}$ by orange points, $\mathbf{v}_{p_{l}}$ and $\mathbf{v}_{p_{r}}$ by orange lines, $\mathbf{v}_{g_{u}}$ by green line, and $\theta^{\prime}$ and $\theta^{\prime\prime}$ by white arc.
@@
2:  Output: Refined mask $M_{r}$// Top View Alignment: First conduct global view alignment to get to ${V^{\prime}}$, then conduct two local view alignments to get ${V^{\prime\prime}}$ and ${V^{\prime\prime\prime}}$, and finally segment the topmost object in ${V^{\prime\prime\prime}}$ to get ${M_{f}}$.
3:  ${V^{\prime}}\leftarrow V$, $V^{\prime\prime}\leftarrow V^{\prime}$, $V^{\prime\prime\prime}\leftarrow V^{\prime\prime}$
4:  ${M_{f}}\leftarrow SAM(V^{\prime\prime\prime}(c))$// Cross-prompted Segmentation: First find four points, then use them to perform segmentation to generate ${M_{s}}$, and finally refine ${M_{s}}$ to obtain ${M_{r}}$ based on ${M_{d}}$.
@@
Through the aforementioned steps, PSP not only effectively segments the topmost object in a densely cluttered scene, which consistently appears near the center of the camera view, but also mitigates the issue of partially visible objects caused by input size constraints of the grasp detection model. It is important to note that this method is not a multi-view grasping [18], which is a complex optimization process. In contrast, our method is more efficient given a single view depth image.
IV-A2 Cross-prompted Segmentation
We propose the cross-prompt to optimize segmentation in PSP because the single-point prompt is highly unstable in dense clutter. This issue is particularly pronounced when the object’s appearance is complex, such as the food packaging, where only part of the object is segmented (usually manifested as many holes in the segmented object). This limitation adversely impacts the subsequent prediction and sampling of grasp candidates.
Specifically, it begins by applying the Sobel operator [31] to extract the edges of the instance mask $M_{f}$ obtained from the initial single-point prompt segmentation. We then search for the two pixels most distant from each other, which we refer to as ${p_{m}}$ and ${p_{m}^{\prime}}$ on the edges. As shown in Eq. 4, $P_{e}$ means the set of all pixels on the edges, and ${P_{e}\times P_{e}}$ represents the Cartesian product of the set $P_{e}$. Next, we calculate the perpendicular line $\perp_{(p_{m},p_{m}^{\prime})}$ connecting ${p_{m}}$ and ${p_{m}^{\prime}}$, and intersecting this perpendicular line with the edges yields another pair of the most distant pixels, which we refer to as ${p_{m_{p}}}$ and ${p_{m_{p}}^{\prime}}$ as shown in Eq. 5 and Fig. 4 (a). These four points are then used as prompts to perform the second segmentation, resulting in ${M_{s}}$. Compared to other points, these two farthest pairs of points can better exploit the geometric constraints of the initial segmentation result, thus recovering the holes in initial segmentation, and then achieving significant results in the second segmentation, as demonstrated in our ablation studies. Finally, ${M_{s}}$ is refined by image dilation processing: a depth threshold is applied and pixels from the first prompt serve as initial points for segmentation to produce ${M_{d}}$. By combining ${M_{d}}$ and ${M_{s}}$, we obtain the refined ${M_{r}}$ as shown in Algorithm 1.
$(p_{m_{p}},p_{m_{p}}^{\prime})=\perp_{(p_{m},p_{m}^{\prime})}\cap{P_{e}}$ (5)
IV-B Monozone Sampling Policy (MSP)
After isolating the topmost object by PSP, we sample the grasp candidates within this object. Here, we discard quality score-based sampling and propose the Monozone Sampling Policy (MSP) to sample grasp candidates for this object. In other words, we do not use the instance mask to guide the quality score-based sampling, but instead focus directly on sampling these grasp candidates within the mask. Moreover, MSP comprehensively considers all parameters of the grasp candidates (rotation $\theta$, position $(x,y)$, and width $w$). It analyzes the relationship between the grasp candidates and the object being grasped and its neighboring objects. Next, we elaborate on these components.
Figure 5: Visualization of adaptive viewpoint rotation in MGA. The red point, green point, and blue point are rotation center $c_{r}$, the center $c_{i}^{\prime}$ of $g_{i}^{\prime}$, and the center $c_{i}$ of $g_{i}$, respectively. The rotation angle is denoted as $\theta_{c}$.
@@ IV-B Monozone Sampling Policy (MSP)
6:  end for// Monozone Grasp Analysis: Filter out $g_{i}$ that leads the robot to collide with this object itself, this object’s adjacent objects, or the table to obtain the optimal grasp $g^{*}$.
@@
IV-B1 Grasp Angle Calibration
Inappropriate selection of the grasp angle $\theta$ may cause the object to slip or fall during grasping due to the uneven force distribution on both fingers of the parallel-jaw gripper. Therefore, all grasp candidates (denoted by $G$) need to be ensured to have an optimal angle before sampling, as it directly determines the subsequent sampling effectiveness. Here, we use the grasping model in [20] to obtain the grasp candidates. Specifically, we first extract the edges of the instance mask. For each grasp candidate, we rotate them clockwise in 2-degree intervals until they reach 360 degrees. For each rotation $r$, we find four intersection points between the two long sides of the grasp candidate and the edges, that is, $p_{t_{l}}$, $p_{b_{l}}$, $p_{t_{r}}$, and $p_{b_{r}}$.
Subsequently, we calculate the angle $\theta^{\prime}$ between the vector $\mathbf{v}_{p_{l}}$ determined by $p_{b_{l}}$ and $p_{t_{l}}$ and the vector $\mathbf{v}_{g_{u}}$ of the long upper side of this grasp candidate, and similarly to get the angle $\theta^{\prime\prime}$ between the vector $\mathbf{v}_{p_{r}}$ determined by $p_{b_{r}}$ and $p_{t_{r}}$ and the vector $\mathbf{v}_{g_{u}}$, shown in Fig. 4 (b). By subtracting 90 degrees from each of these angles, taking the absolute value, and summing them, we obtain the angle difference for each rotation. Finally, we select the rotation with the smallest angle difference ${R}^{*}$ given by Eq. 6 and use this rotated box as the new grasp candidate.
[…]
IV-B2 Monozone Grasp Analysis
After grasp angle calibration, we individually analyze each grasp candidate $g_{i}$ to determine the optimal grasp. Let $G^{\prime}$ denote the grasp candidate sets after angle calibration. We first examine the relationship between each grasp candidate $g_{i}$ and the instance mask $M_{r}$ by checking whether the pixels ${P_{i}}^{s}$ along the two short sides of the grasp candidate fall within the instance mask. We filter out $g_{i}$ if any pixel $p$ in ${P_{i}}^{s}$ is within mask $M_{r}$. Additionally, to ensure that grasp candidates are concentrated within $M_{r}$, we also filter out $g_{i}$ whose centers $c_{i}$ are not within $M_{r}$ and get grasp candidate sets $G^{\prime\prime}$, which is shown in Eq. 7.
However, since these processes probably filter all $g_{i}$ out, we adaptively rotate the viewpoint clockwise to alter the grasp candidate. That is, if no $g_{i}$ are available from the current viewpoint, we rotate the image 30 degrees at a time and repeat it until the available grasp candidate is found. Since the camera does not rotate and is constrained by hand-eye calibration, we need to project the rotated candidate grasp $g_{i}^{\prime}$ back to the original viewpoint. Here, the parameters $w^{\prime}$ and $h^{\prime}$ of $g_{i}^{\prime}$ remain unchanged. The angle $\theta^{\prime}$ can be adjusted by adding the rotation angle $\theta_{c}$ and restricting it to the range $[-\pi/2,\pi/2]$ to convert it back to the angle $\theta$ of the candidate grasp $g_{i}$ under the original viewpoint. For the center $c_{i}^{\prime}(x^{\prime},y^{\prime})$ of $g_{i}^{\prime}$, assuming the center of rotation is $c_{r}(x_{r},y_{r})$, the projection relationship from $c_{i}^{\prime}(x^{\prime},y^{\prime})$ to the center $c_{i}(x,y)$ of the grasp candidate $g_{i}$ under the original viewpoint can be obtained from Eq. 8. The visualization of viewpoint rotation is shown in Fig. 5.
$\begin{bmatrix}x\\ y\end{bmatrix}=\begin{bmatrix}\cos(\theta_{c})&-\sin(\theta_{c})\\ \sin(\theta_{c})&\cos(\theta_{c})\end{bmatrix}\begin{bmatrix}x^{\prime}-x_{r}\\ y^{\prime}-y_{r}\end{bmatrix}+\begin{bmatrix}x_{r}\\ y_{r}\end{bmatrix}$ (8)
Next, based on $G^{\prime\prime}$, we analyze their relationship with adjacent objects by setting a depth threshold $T_{d}$, that is, if the depth difference between any $p$ and $c_{i}$ exceeds $T_{d}$, the grasp candidate $g_{i}$ will be filtered out and get grasp candidate sets $G^{\prime\prime\prime}$, as shown in Eq. 9. The above steps can filter out most grasp candidates $g_{i}$ that could lead the robot to collide with a target object or its adjacent objects, as well as the table.
Figure 6: Visualization of OCID grasping dataset. Each subfigure shows different piled objects in different backgrounds.
Figure 7: Objects for the grasping experiment: toys, ragdolls, household goods, and snacks (clockwise from top left).
Finally, based on the sampled grasp candidate sets $G^{\prime\prime\prime}$, we use our previous method [32] to select the $g_{i}$ with the smallest center pixel depth value $d(c_{i})$ as the optimal grasp $g^{*}$, which is shown in Eq. 10. This operation ensures a more stable and safer grasp, especially when dealing with objects with complex shapes.
IV-B3 Optimal Grasp Refinement
Although the optimal grasp $g^{*}$ was obtained through the previous coarse-to-fine analysis, this analysis will be influenced by random errors from the depth camera, where $g^{*}$ might still result in collisions with adjacent objects during grasping execution. One way to get around this problem was reported in [33], where a series of intervals was defined within the grasp box and the grasp width and position were adjusted based on the relationships between these intervals. However, this method relies on local regions of the depth image with errors, and it is computationally cumbersome. Therefore, we find the minimum rectangle $R_{ec}$ intersecting the optimal grasp $g^{*}$ and the instance mask $M_{r}$, and this process can be represented by $f(g^{*},M_{r})$. Then calculate the shortest width $w_{s}$ and a new center point $c^{\prime}_{i}$ of this grasp through ${min{R_{ec}}(w)}$ and ${R_{ec}}(c)$. Additionally, to mitigate the impact of hand-eye calibration errors, we further expand $w_{s}$ to $w^{\prime}_{s}$ by adding some of the hand-eye calibration translation errors $e_{c}$ in the $X$ and $Y$-axis. Finally, we use $w^{\prime}_{s}$ and $c^{\prime}_{i}$ as the new width and center of the grasp for optimal grasp, and denote the final grasp by $g^{*}_{f}$, as shown in Eq. 11, Eq. 12, and Algorithm 2.
$R_{ec}=f(g^{*},M_{r})$ (11)
@@ V Experiments
In this section, we validate the effectiveness of our method by conducting benchmarking studies. Firstly, we compare it with baseline grasping methods in various mid-clutter (up to 20 objects) and high-clutter (up to 50 objects) scenes. Then we increase the number of objects to 100 (extreme-clutter) and analyze the effectiveness of PSP and MSP.
@@ V-A1 Setting for Grasping Model
The baseline methods are categorized into two groups. The first group includes GGCNN [17], GGCNN2 [18], GRconvnet [20], SEnet [19], and FCGnet [21], which are suitable for mid-clutter scenarios. The second group comprises DexNet 4.0 [4] and GraspNet [5], which are tailored for high-clutter scenarios.
[…]
@@
Figure 8: Line graph showing GSR of PMSGP and first-group baselines. The horizontal axis represents different methods and the depth axis represents trials from T1 to T5. The vertical axis represents the number of grasp failures. We emphasize the number of grasp failures (T1, T3, T5) in each method with dots, and connect them with dashed lines to better show the difference.
TABLE I: GSR comparison among PMSGP and first-group baselines
@@
Figure 9: Bar graphs showing GSR of PMSGP and second-group baselines. (a), (b), (c), and (d) represent the results of testing ragdolls, snacks, toys, and household goods. In each subfigure, the vertical axis represents the number of grasp failures, and the horizontal axis represents different methods with five trials. We show the positive and negative errors at the top of each bar by calculating the mean of the number of grasp failures across all trials for each method.
TABLE II: GSR comparison among PMSGP and second-group baselines
Methods Ragdolls Snacks Toys Household goods
@@
Figure 10: Performance of PMSGP on objects of different grasping difficulties and scene complexities. The first row is a mid-clutter scene consisting of 20 ragdolls, the second row is a high-clutter scene consisting of 50 snacks, and the last row is an extreme-clutter scene consisting of 100 household goods. The topmost object in each subfigure is emphasized by green lines.
V-A2 Setting for Real Grasping
Our grasping system consists primarily of an Intel RealSense D435 depth camera and a UFactory xArm 5 robot. We employ an eye-in-hand architecture, with the camera mounted on the robot’s distal end and facing downward. The objects used in our grasping experiments are divided into four categories, with a total of 300 novel objects: 50 ragdolls (Category 1), 100 snacks (Category 2), 50 toys (Category 3), and 100 household goods (Category 4), respectively, as shown in Fig. 7. Category 1 is the easiest and as the category level goes up, the grasping difficulty rises, too.
Before grasping, we fill the depth hole and set a depth value threshold (with the upper limit of 40 $cm$ and the lower limit of 10 $cm$ explained earlier in Section IV-A) to ensure that the grasp is executed within this range [17]. During grasping, each method is tested in five trials per experiment, and the number of failed grasps in each trial (T) is recorded. The grasp success rate (GSR) is calculated by dividing the total number of successful grasps by the total number of grasp attempts across five trials. In addition, to improve experimental safety and ensure all objects are grasped in each trial, we provide minimal manual assistance during the experiments. Specifically, if an object fails to be grasped 2-3 times, we manually pick up the object and count it as a failure. Additionally, if an object moves out of the camera view, it is repositioned with manual intervention. Similarly, if a grasped object moves out of the robot’s workspace, causing it to stop, the object is repositioned manually, too.
V-B Comparision Studies
V-B1 Comparision with Baseline Methods in Mid-clutter
We compare PMSGP with the baseline methods in the first group. We used 10 snacks and 10 household goods to form a mid-clutter scene. The results are shown in Table I. PMSGP achieves a GSR of 93.5% (100/107), with only 7 grasp failures, which is far superior to other baselines except for FCGnet. Additionally, we found that some baselines perform well on the dataset but not in real grasping. For example, GGCNN2 has a GSR of only 47.6% (100/210) with a total of 110 grasp failures, indicating that this method does not generalize well to novel objects in mid-clutter. Finally, we also visualize the result in Fig. 8 to better show the gap between PMSGP with other baseline methods. As shown in this figure, it is very clear that our method’s number of failures is much smaller with little variance across the trials.